OpenStudy (anonymous):

find the exact values of sin2u, cos2u, tan2u using cotu = sqrt2, pi < u < 3pi/2

OpenStudy (jdoe0001):

$$ cot(u) = \sqrt{2} \implies cot(u) = \cfrac{cos(u)}{sin(u)}\\ \text{for the section between } \pi < u < \cfrac{3\pi}{2} \implies 3rd \ quadrant\\ \text{sine and cosine are both negative on 3rd quadrant, so}\\ cot(u) = \cfrac{cos(u)}{sin(u)} = \cfrac{-\sqrt{2}}{-1}\\ $$

OpenStudy (jdoe0001):

so, that'd give you a cosine of \(-\sqrt{2}\) and a sine of -1 from there you can use that to expand the "double angle trig identities" and get their value

OpenStudy (dumbcow):

ok you are given angle is in 3rd quadrant (sin is neg, cos is neg) start with triangle, then you will need the double angle formulas  from the triangle you obtain \[\sin u = \frac{1}{\sqrt{3}}\] \[\cos u = \frac{\sqrt{2}}{\sqrt{3}}\] \[\tan u = \frac{1}{\sqrt{2}}\]

OpenStudy (dumbcow):

oops i forgot about neg signs

OpenStudy (dumbcow):

@jdoe0001 , careful with assigning cos and sine , they have to be within -1 and 1 sqrt2 > 1

OpenStudy (jdoe0001):

http://www.sosmath.com/trig/Trig5/trig5/img7.gif

OpenStudy (jdoe0001):

oohh yeah, my bad $$ cot(u) = \sqrt{2} \implies cot(u) = \cfrac{adjacent}{opposite}\\ \text{for the section between } \pi < u < \cfrac{3\pi}{2} \implies 3rd \ quadrant\\ \text{"y" and "x" are both negative on 3rd quadrant, so}\\ cot(u) = \cfrac{adjacent}{opposite} = \cfrac{-\sqrt{2}}{-1}\\ $$

OpenStudy (anonymous):

So cosine is negative in Quadrant 3? But sqrt2 > 1, so that means that cosine can't equal - sqrt2.

OpenStudy (jdoe0001):

yes, as properly indicated by dumbcow

OpenStudy (jdoe0001):

so, we're just dealing with adj/opp sides for this case

OpenStudy (anonymous):

\[\cot u=\sqrt{2},\tan u=\frac{ 1 }{ \sqrt{2} }\] \[\sin 2u=\frac{ 2\tan u }{ 1+\tan ^{2}u }=\frac{ 2*\frac{ 1 }{ \sqrt{2} } }{1-\frac{ 1 }{ 2 } }\] \[=\frac{ \frac{ 2 }{\sqrt{2}} }{\frac{ 1 }{2 } }=2\sqrt{2}\] \[similarly \cos 2u=\frac{ 1-\tan ^{2}u }{1+\tan ^{2}u } ,\tan 2 u=\frac{ 2\tan u }{ 1-\tan ^{2}u }\] Plug the values of tan u ,you get the required values.

OpenStudy (anonymous):

I used 2sinucosu for sin2u and got 2sqrt2/3.

OpenStudy (anonymous):

How did you go from 1 + tan^2u to 1 - (1/2). Isn't is 1 + (1/2)?

OpenStudy (dumbcow):

@rascal , you are correct , that was a typo i believe either formula will work, however 2sincos is much more familiar

OpenStudy (anonymous):

Ok thanks!

OpenStudy (anonymous):

\[\sin 2u=\frac{ 2\tan u }{ 1+\tan ^{2}u }\] \[\tan 2u=\frac{ \sin 2u }{ \cos 2u }\] \[\cos 2u=\frac{ 1-\tan ^{2}u }{1+\tan ^{2}u }\]